OpenStudy (anonymous):

Is the system consistent or inconsistent? y= -1/2x-1 and y=3x+3

OpenStudy (anonymous):

i am working on it

OpenStudy (anonymous):

r u still there?

OpenStudy (anonymous):

Yes, haha.

OpenStudy (anonymous):

i am working on it

OpenStudy (anonymous):

The system of equations a1x + b1y + c1 = 0 a2x + b2y + c2 = 0 (i) is consistent with unique solution, if a1/a2 is not equal to b1/b2 i.e the lines represented by equations are not parallel. (ii) consistent with infinitely many solutions, if a1/a2 = b1/b2 = c1/c2 i.e. the lines represented by equations are coincident. (iii) inconsistent, if a1/a2 = b1/b2 is not equal to c1/c2 i.e. the lines represented by equations are parallel and non-coincident. Now use this result and find the answer yourself :)